# WallFacer: Guiding Transformer Model Training Out of the Long-Context Dark Forest with N-body Problem

### v2

@@ 5. Evaluation
Figure 8. The normalized relative memory cost of different configurations of WallFacer compared with Ring Attention on different clusters.
5.1. Throughput and Adaptability
@@
The results of our evaluation are illustrated in Figure 7. We measure throughput in thousands of tokens per second. To better demonstrate how to select the optimal configuration of WallFacer under each condition, we included two configurations, Wall-2 and Wall-4, in the figure. We omit configurations with lower performance for clarity. As indicated in the figure, in all six settings, at least one configuration of WallFacer achieves higher throughput than Ring Attention, with performance improvements of 62.87%, 42.45%, 35.98%, 19.9%, 77.12%, and 34.62%. This advantage is primarily due to the additional parallel dimension that WallFacer introduces. Unlike Ring Attention, which requires inter-node P2P communication in each iteration, WallFacer’s P2P communication is mostly confined intra-node, except for initial data transfers. This experiment clearly demonstrates WallFacer’s superior performance across various environments.
@@
5.2. Memory Consumption
@@
Memory consumption in our experiments stems from both the model weights and activations, with additional costs for WallFacer arising from the duplication of QKV matrices prior to attention computation. To quantitatively assess the additional memory required by WallFacer, we monitored the maximum memory allocated by PyTorch (Paszke et al., 2019) during our experiments. It is important to note that memory fragmentation in PyTorch can impact memory allocation efficiency. To mitigate this, we limited the sequence lengths in our training to prevent PyTorch from triggering cuda_free when the allocated memory approaches the GPU’s limit, which would otherwise introduce significant overhead.
@@
The results, displayed in Figure 8, reveal that for the configurations yielding the highest throughput, Ring Attention consumes between 7.9% and 30.79% less GPU memory than WallFacer. However, considering the substantial throughput gains provided by WallFacer, the additional memory usage is deemed acceptable. Additionally, evaluations for long sequences typically employ sequence lengths that are powers of two, and we observed that the maximum supported sequence lengths were not significantly impacted by this additional memory usage.
@@
Moreover, in scenarios involving larger models with thirty billion (30B) or seventy billion (70B) parameters, the relative increase in memory consumption due to QKV duplication diminishes. This reduction is explained by equation 10, which shows that as the model size increases—owing to more parameters and additional Transformer layers—the proportion of memory consumed by model weights and activations grows, whereas the absolute extra memory incurred by QKV multiplication remains constant. This phenomenon is further evidenced by the fact that the extra memory ratio for experiments with the 7B model is significantly smaller than that for the 3B model.
Figure 9. Strong scaling experiments on Nvidia H100 GPUs with fixed sequence length of 128K.
5.3. Strong and Weak Scaling
@@
In the scaling tests we carry out experiments for both strong and weak scaling. Strong Scaling maintains the scale of the problem that we are trying to solve while increasing the computing resource that we use to speed it up. So in this experiment, we fix the sequence length to 128K while increasing the number of GPUs from 8 to 64. As id depicted in Figure 9, WallFacer shows gradually more advantage over Ring Attention as we increase the number of GPUs. When we use 64 GPUs, WallFacer shows 60.71% higher throughput than Ring Attention. This can also explained by the computation-communication ratio. When scaled to more GPUs, the local sequence length on each GPU becomes smaller, and as explained the previous sections, makes it harder to overlap the P2P communication with attention computation. This is exactly the main advantage of WallFacer, reduction of total P2P communication volume. So WallFacer can give a promising performance when we are trying to speedup the training of a fixed amount of data with more GPUs.
Figure 10. Weak scaling on Nvidia H100 GPUs of sequence length from 64K to 512K
As for weak scaling, we can see from Figure 10 that WallFacer can still show gradually more advantage over Ring Attention, but not as much as that in strong scaling experiments. This is because the computation-communication ratio stays the same in weak scaling. But with more GPUs, Ring Attention will require a higher portion of inter-node communication, while WallFacer can remain most communication local.
@@
6. Related Works
@@
Attention Optimization. Traditional full attention mechanisms necessitate $O(n^{2})$ memory for storing the outputs of $QK^{T}$, leading to significant computational and memory demands. To address these challenges within the GPU, several approaches have been devised to reduce both memory and computational requirements. Memory-efficient attention(Rabe and Staats, 2022) introduces a straightforward algorithm that requires only $O(1)$ memory relative to the sequence length, with an extension for self-attention that needs only $O(\log n)$ memory. Flash Attention further minimizes I/O overhead and enhances overall efficiency. Additionally, optimization methods specifically tailored for inference, such as PagedAttention(Kwon et al., 2023), are also being developed to improve the efficiency of attention computations. In this work, we utilize Flash Attention within each iteration to reduce the computation overhead.
@@
Long-Sequence Training Techniques. In recent years, numerous techniques have been developed for long-sequence training. Sequence Parallelism(Li et al., 2021) was initially introduced to enhance the efficiency of parallel long-sequence training. Ring Attention(Liu et al., 2023b) improved communication efficiency through memory-efficient methods(Rabe and Staats, 2022), supporting near-infinite sequence lengths. DeepSpeed Ulysses(Jacobs et al., 2023) employs attention head splitting to achieve high efficiency, though it is constrained by the number of heads. Megatron Sequence Parallelism focuses on reducing memory costs during Tensor Parallelism, while DistFlashAttention(Li et al., 2024) features a load-balance scheme and a novel gradient checkpoint method. Our work builds on these innovations, introducing a system that supports near-infinite sequence length and large-scale training with an efficient communication scheme.
@@
Techniques for Distributed Model Training. Distributed model training encompasses two primary areas: 1) Memory Management: Various techniques aim to conserve GPU memory during distributed training, such as mixed precision training(Micikevicius et al., 2018) and the ZeRO series(Rajbhandari et al., 2020). Additionally, Zero-Offload(Ren et al., 2021) leverages CPU memory offloading. In this work, we implement ZeRO-2 to manage optimizer states and gradients efficiently. 2) Hybrid Parallelism: Frameworks like Megatron(Narayanan et al., 2021) and Colossal AI(Bian et al., 2021) integrate multiple forms of parallelism. There are various existing Parallelism techniques like Pipeline Parallelism(Huang et al., 2018; Fan et al., 2020; Li and Hoefler, 2021; Liu et al., 2023a) and Tensor Parallelism(Shoeybi et al., 2019), which can be combined with WallFacer Parallelism to facilitate large-scale training. We are also considering the integration of additional frameworks such as (Chen et al., 2024) to enhance overlapping capabilities in future implementations. In the future, we will examine hybrid strategies that combine WallFacer with other parallelisms.
7. Conclusion
@@
WallFacer represents an advanced near-infinite-context Transformer model training system, featuring a communication-optimized multi-ring sequence parallelism scheme. In this study, we also conceptualize attention computation as a special case of the n-body problem, which introduces new perspectives and potential solutions for future challenges. Through comprehensive adaptability and scaling experiments, we demonstrate that our system not only achieves high efficiency across various training environments but also excels under both strong and weak scaling conditions. In an era increasingly demanding longer contexts for both natural language processing and computer vision, WallFacer is poised to make significant contributions to the industry and inspire innovative research in academia.

### v3

@@ 4 Evaluation
Figure 8: The normalized relative memory cost of different configurations of WallFacer compared with Ring Attention on different clusters.
4.1 Throughput and Adaptability
@@
The results of our evaluation are illustrated in Figure 7. We measure throughput in thousands of tokens per second. To better demonstrate how to select the optimal configuration of WallFacer under each condition, we included two configurations, Wall-2 and Wall-4, in the figure. We omit configurations with lower performance for clarity. As indicated in the figure, in all six settings, at least one configuration of WallFacer achieves higher throughput than Ring Attention, with performance improvements of 114.33%, 41.4%, 62.87%, 42.45%, 35.98%, 19.9%, 77.12%, and 34.62%. This advantage is primarily due to the additional parallel dimension that WallFacer introduces. Unlike Ring Attention, which requires inter-node P2P communication in each iteration, WallFacer’s P2P communication is mostly confined intra-node, except for initial data transfers. This experiment clearly demonstrates WallFacer’s superior performance across various environments.
@@
Table 4: Supported Sequence Length of Ring Attention and WallFacer on one Nvidia A100 80GB GPU.
Supported Seq Len on one 80GB A100 GPU
(K Tokens)
Model Size Length Ring Attention WallFacer
3B 128
256
512
7B 128
256
512
13B 128
256
512
4.2 Memory Consumption
Before evaluating the memory consumption of WallFacer, we first compared the supported sequence lengths of WallFacer with those reported in the RingAttention paper [liu2023ring]. As shown in Table 4, despite the slightly higher memory requirements of WallFacer, it still supports sequence lengths commonly used in training tasks.
@@ 4.2 Memory Consumption
Memory consumption in our experiments stems from both the model weights and activations, with additional costs for WallFacer arising from the duplication of QKV matrices prior to attention computation. To quantitatively assess the additional memory required by WallFacer, we monitored the maximum memory allocated by PyTorch [pytorch] during our experiments. It is important to note that memory fragmentation in PyTorch can impact memory allocation efficiency. To mitigate this, we limited the sequence lengths in our training to prevent PyTorch from triggering cuda_free when the allocated memory approaches the GPU’s limit, which would otherwise introduce significant overhead.
@@
The results, displayed in Figure 8, reveal that for the configurations yielding the highest throughput, Ring Attention consumes between 7.9% and 30.79% less GPU memory than WallFacer, except the DiT 512k Wall-4 setting. However, considering the substantial throughput gains provided by WallFacer, the additional memory usage is deemed acceptable. We observed that the maximum supported sequence lengths were not significantly impacted by this additional memory usage.
@@
Moreover, in scenarios involving larger models, the relative increase in memory consumption due to QKV duplication diminishes. This reduction is explained by equation 7, which shows that as the model size increases—owing to more parameters and additional Transformer layers—the proportion of memory consumed by model weights and activations grows, whereas the absolute extra memory incurred by QKV multiplication remains constant. This phenomenon is further evidenced by the fact that the extra memory ratio for experiments with the 7B model is significantly smaller than that for the 3B and 1B model.
4.3 Strong and Weak Scaling
@@
In the scaling tests we carry out experiments for both strong and weak scaling. Strong Scaling maintains the scale of the problem that we are trying to solve while increasing the computing resource that we use to speed it up. So in this experiment, we fix the sequence length to 128K while increasing the number of GPUs from 8 to 64 for the GPT model and from 8 to 32 for the DiT model. As id depicted in Figure LABEL:Fig.strong, WallFacer shows gradually more advantage over Ring Attention as we increase the number of GPUs. When we use 64 GPUs, WallFacer shows 60.71% higher throughput than Ring Attention for the GPT model and 114.33% for the DiT model on 32 A100 GPUs. This can also explained by the computation-communication ratio. When scaled to more GPUs, the local sequence length on each GPU becomes smaller, and as explained the previous sections, makes it harder to overlap the P2P communication with attention computation. This is exactly the main advantage of WallFacer, reduction of total P2P communication volume. So WallFacer can give a promising performance when we are trying to speedup the training of a fixed amount of data with more GPUs.
In our weak scaling experiments using DiT models, we scale the sequence length from 128k to 512k while proportionally increasing the number of GPUs from 8 to 32. Assuming a scaling factor of $k$, both the sequence length and the number of GPUs increase by $k$, resulting in a total computation volume that increases by approximately $k^{2}$. Since the computational resources also increase by $k$, the theoretical overall throughput (tokens per second) should remain constant during this weak scaling setting. As shown in Figure LABEL:Fig.weak, both WallFacer and Ring Attention exhibit this expected behavior, with throughput remaining relatively constant. However, WallFacer consistently achieves higher throughput across different numbers of GPUs. This is because, although the local computation-to-communication ratio stays the same in weak scaling, Ring Attention requires a higher proportion of inter-node communication as the number of GPUs increases, whereas WallFacer keeps most communication local.
@@
5 Discussion
5.1 FlashAttention3 and Hopper GPUs
In addition to the original FlashAttention [dao2022flashattention] used in our experiments, FlashAttention3 [shah2024flashattention3fastaccurateattention] has been introduced, specifically designed for Hopper and newer Nvidia GPUs. For FP16 precision, which is utilized in this paper, FlashAttention3 achieves a 1.5-2.0x speedup on Hopper GPUs. As discussed in Section 2, reducing attention computation overhead results in more P2P communication not being overlapped, further emphasizing the need to reduce communication volume. With the increasing adoption of Hopper GPUs, the significance of the WallFacer system will also grow.
5.2 WallFacer and Other Parallelisms
Model Parallelism As is well known, tensor parallelism shards activations by attention heads during attention computation, making it easily combinable with WallFacer with minimal effort. However, when combined with tensor parallelism, the need for attention heads can limit the scalability of head-based sequence parallel methods like Ulysses. Pipeline parallelism, on the other hand, divides the model across layers without altering the computation patterns within Transformer blocks, making WallFacer orthogonal to it.
Other Sequence Parallelism WallFacer is orthogonal with other attention-head-sharding-based sequence parallelism approaches, such as DeepSpeed Ulysses [jacobs2023Ulysses]. While Ulysses distributes attention heads across different devices, WallFacer can independently partition activations along the sequence length dimension. In future work, we can explore combining WallFacer with DeepSpeed Ulysses to expand the communication scheduling space, harnessing the scalability of WallFacer alongside the efficiency of Ulysses.
In summary, WallFacer can be seamlessly integrated with other parallel training techniques, enabling the creation of a hybrid distributed training system.
6 Related Work
@@
Attention Optimization. Traditional full attention mechanisms necessitate $O(n^{2})$ memory for storing the outputs of $QK^{T}$, leading to significant computational and memory demands. To address these challenges within the GPU, several approaches have been devised to reduce both memory and computational requirements. Memory-efficient attention[rabe2022selfattention] introduces a straightforward algorithm that requires only $O(1)$ memory relative to the sequence length, with an extension for self-attention that needs only $O(\log n)$ memory. Flash Attention further minimizes I/O overhead and enhances overall efficiency. Additionally, optimization methods specifically tailored for inference, such as PagedAttention[kwon2023efficient], are also being developed to improve the efficiency of attention computations. In this work, we utilize Flash Attention within each iteration to reduce the computation overhead.
@@
Long-Sequence Training Techniques. Sequence Parallelism[Li2021SequencePL] was initially introduced to enhance the efficiency of parallel long-sequence training. Ring Attention[liu2023ring] improved communication efficiency through memory-efficient methods[rabe2022selfattention], supporting near-infinite sequence lengths. DeepSpeed Ulysses[jacobs2023Ulysses] employs attention head splitting to achieve high efficiency, though it is constrained by the number of heads. Megatron Sequence Parallelism focuses on reducing memory costs during Tensor Parallelism, while DistFlashAttention[li2024distflashattn] features a load-balance scheme and a novel gradient checkpoint method. Our work builds on these innovations, introducing a system that supports large-scale training with an efficient communication scheme.
@@
Techniques for Distributed Model Training. Distributed model training encompasses two primary areas: 1) Memory Management: Various techniques aim to conserve GPU memory during distributed training, such as mixed precision training[micikevicius2018mixed] and the ZeRO series[rajbhandari2020zero]. In this work, we implement ZeRO-2 to manage optimizer states and gradients efficiently. 2) Hybrid Parallelism: Frameworks like Megatron[megatron] and Colossal AI[bian2021colossal] integrate multiple forms of parallelism. There are various existing Parallelism techniques like Pipeline Parallelism[gpipe, dapple, li2021chimera, Liu2023Hanayo] and Tensor Parallelism[shoeybi2019megatron], which can be combined with WallFacer Parallelism to facilitate large-scale training. We are also considering the integration of additional frameworks such as [chang2024centauri] to enhance overlapping capabilities in future implementations.
7 Conclusion
@@
WallFacer represents an advanced near-infinite-context Transformer model training system, featuring a communication-optimized multi-ring sequence parallelism scheme. Through comprehensive adaptability and scaling experiments, we demonstrate that our system not only achieves high efficiency across various training environments but also excels under both strong and weak scaling conditions for both Computer Vision(CV) and Natral Language Processing(NLP) models. In an era increasingly demanding longer contexts for both NLP and CV, WallFacer is poised to make significant contributions to the industry and inspire innovative research in academia.
[…]